Mr.glidden is expanding the size of his garden. His new garden will be 3 times as long and 2 times as wide as his original garden. The perimeter of his original garden is 30 feet and the length is 11 feet.What will be the perimeter in feet of his new garden?

Answer by lwsshak3(11628)

Rectangles/1003722 (2015-11-14 21:40:44): Mr.glidden is expanding the size of his garden. His new garden will be 3 times as long and 2 times as wide as his original garden. The perimeter of his original garden is 30 feet and the length is 11 feet.What will be the perimeter in feet of his new garden?
let x= width of original garden
2*11+2x=perimeter of original garden=30
2x=30-22=8
x=4
2x=8=width of new garden
3*11=33=length of new garden
2*8*+2*33=16+66=82 ft=perimeter of new garden